Find the values of x and y that solve the following system of equations
-5x-7y=-1
2x+9y=-12

Answer by ptaylor(2198)

-5x-7y=-1 ---------------------eq1
2x+9y=-12----------------------eq2

multiply each term in eq1 by 2 and each term in eq2 by 5, then add the two equations:
-10x-14y=-2--------------------------eq1a
10x+45y=-60------------------------------eq2a
adding the two equations, we get:
31y=-62 divide both sides by 31
y=-2---------------------------------------y value
substitute y=-2 into eq2
2x+9(-2)=-12
2x-18=-12 add 18 to each sid
2x-18+18=-12+18 collect like terms
2x=6 divide each side by 2
x=3-----------------------------------x value
CK
substitute the x and y values into eq1:
(-5)(3)-(7)(-2)=-1
-15+14=-1
-1=-1
substitute the x and y values into eq2:
2(3)+9(-2)=-12
6-18=-12
-12=-12


Hope this helps---ptaylor